# ZTF-EP-24-09

### v1

@@ 4 Layered crust
Even if we set aside corrections stemming from anisotropy and asphericity, Runcorn’s model subtly relies on one more crucial assumption, “instant” acquisition of the magnetization throughout the whole crust. A more realistic scenario was analysed by Srnka [Srn76a], in which the thermoremanent magnetization is acquired as the astrophysical body cools inwards and magnetic field of each magnetized layer, in addition to the core’s magnetic field, magnetizes newly cooled layers. Additional permeable effects were treated [^1] This line of research builds on previous Stephenson’s analysis [Ste76b, Ste76a]. In the Stephenson’s model all layers, each with different permeability, are initially embedded in the core’s magnetic field, then they instantly acquire thermoremanent magnetization proportional to the ambient magnetic field, after which the core’s magnetic field is “turned-off”. in the accompanied proceedings paper [Srn76b], by taking into account permeability of the layers which were previously magnetized.
@@
Magnetic field corresponding to this magnetized shell is, as a consequence of the generalized Runcorn’s theorem, zero in the external region $r>r_{0}$, but nonzero in the interior region $r<r_{1}$ (details are derived in the Appendix),
@@
Major obstacle here is the (lack of) knowledge about the average layer thickness $d$, as well as the constants $x_{m}$ and $x_{M}$ (in the lunar case Srnka argues for the estimate $d\sim 1\,$km). A simple shortcut is to use Taylor series $1-(1+\varepsilon)^{-n}=n\varepsilon+O(\varepsilon^{2})$ and rough estimates $x_{m}\sim d/r_{0}$ and $x_{M}\sim d/r_{N}$, in order to produce estimates for the lower bound $(2\ell+1)(1-r_{N}/r_{0})$ and the upper bound $(2\ell+1)(r_{0}/r_{N}-1)$. However, these values must be taken with a grain of salt, as any more credible bounds require further information about the interior structure of the crust or generalization of the model for the continuous cooling.
@@ 5 Final remarks
Runcorn’s theorem is one of the hidden gems of magnetostatics, which has proven as an important insight about the magnetic field of magnetized crusts of astrophysical bodies. In section 2 we have offered one possible bird’s-eye view on the mathematical origin of the Runcorn’s result. Focal results here is the theorem 2.1, revealing specific orthogonality of gradients of harmonic functions on the spherical shell in the intersection of their domains. An immediate corollary is the vanishing of the exterior (interior) magnetic field for the spherical shell magnetized by the interior (external) source. In section 3 we have analysed several deviations of the Runcorn’s theorem, coming from anisotropic susceptibility and/or nonsphericity of the crust. Inequalities (21) and (24) quantify an intuitive expectation that the magnetic field in these cases is bounded by the dipole, $O(r^{-3})$ term. Finally, in section 4 we have revisited a model of a crust magnetized during the inward cooling, with the sharpening of the Srnka’s estimates.
@@ Acknowledgements
An elementary problem that was used in the section 4 consists of a single spherical shell $r_{-}<r<r_{+}$ with magnetization $\bm{M}$, surrounded by vacuum. Magnetic scalar potential, introduced via $\bm{H}=-\nabla\Psi$, is a harmonic function, which in the spherical coordinate system may be written in the form
@@
Coefficients $B_{\ell m}^{\mathrm{(ext)}}$, $A_{\ell m}$, $B_{\ell m}$ and $A_{\ell m}^{\mathrm{(in)}}$ may be found from the junction conditions, continuity of the potential $\Psi$ and continuity of the normal component of the magnetic field $\bm{B}=\mu_{0}(\bm{H}+\bm{M})$ on spheres $r=r_{-}$ and $r=r_{+}$. We shall consider two separate subcases:
@@
    Magnetization of the form $\bm{M}=-\nabla\sum_{\ell,m}M_{\ell m}r^{-(\ell+1)}\mathrm{Y}_{\ell m}$. Solution of junction conditions are
@@
    Magnetization of the form $\bm{M}=-\nabla\sum_{\ell,m}M_{\ell m}r^{\ell}\mathrm{Y}_{\ell m}$. Solution of junction conditions are
@@
  • [LL01] E. H. Lieb and M. Loss, Analysis, 2nd ed. ed., American Mathematical Society, Providence, 2001.

### v2

@@ 4 Layered crust
Even if we set aside corrections stemming from anisotropy and asphericity, Runcorn’s model subtly relies on one more crucial assumption, “instant” acquisition of the magnetization throughout the whole crust. A more realistic scenario was analysed by Srnka [Srn76a], in which the thermoremanent magnetization is acquired as the astrophysical body cools inwards and magnetic field of each magnetized layer, in addition to the core’s magnetic field, magnetizes newly cooled layers. Additional permeable effects were treated [^2] This line of research builds on previous Stephenson’s analysis [Ste76b, Ste76a]. In the Stephenson’s model, all layers, each with different permeability, are initially embedded in the core’s magnetic field, and then they instantly acquire thermoremanent magnetization proportional to the ambient magnetic field, after which the core’s magnetic field is “turned-off”. in the accompanied proceedings paper [Srn76b], by taking into account permeability of the layers which were previously magnetized.
@@
Magnetic field corresponding to this magnetized shell is, as a consequence of the generalized Runcorn’s theorem, zero in the external region $r>r_{0}$, but nonzero in the interior region $r<r_{1}$ (details are derived in Appendix),
@@
Major obstacle here is the (lack of) knowledge about the average layer thickness $d$, as well as the constants $x_{m}$ and $x_{M}$ (in the lunar case Srnka argues for the estimate $d\sim 1\,$km). A simple shortcut is to use Taylor series $1-(1+\varepsilon)^{-n}=n\varepsilon+O(\varepsilon^{2})$ and rough estimates $x_{m}\sim d/r_{0}$ and $x_{M}\sim d/r_{N}$, in order to produce estimates for the lower bound $(2\ell+1)(1-r_{N}/r_{0})$ and the upper bound $(2\ell+1)(r_{0}/r_{N}-1)$. These values, however, must be taken with a grain of salt, as any more credible bounds require further information about the interior structure of the crust.
@@ 5 Final remarks
Runcorn’s theorem is a hidden gem of magnetostatics, providing an important insight into the magnetic field of magnetized crusts of astrophysical bodies. In Section 2 we have offered one possible bird’s-eye view on the mathematical origin of the Runcorn’s result. Focal result here is Theorem 2.1, revealing specific orthogonality of gradients of harmonic functions on the spherical shell in the intersection of their domains. An immediate corollary is the vanishing of the exterior (interior) magnetic field for the spherical shell magnetized by the interior (external) source.
[…]
Finally, in Section 4 we have revisited a model of a crust magnetized during the inward cooling, with the sharpening of the Srnka’s estimates. It would be interesting to find a generalization of the latter model, to a crust which is magnetized during the continuous process of inward cooling. Also, one could take into account various external magnetic fields, mentioned in the discussion of the generalized Runcorn’s Theorem 2.1, and study their effect on the crustal magnetization.
Data availability statement
No new data were created or analysed in this study. Data sharing is not applicable to this article.
@@ Acknowledgements
An elementary problem that was used in Section 4 consists of a single spherical shell $r_{-}<r<r_{+}$ with magnetization $\bm{M}$, surrounded by vacuum. Magnetic scalar potential, introduced via $\bm{H}=-\nabla\Psi$, is a harmonic function, which in the spherical coordinate system may be written in the form
@@
Coefficients $B_{\ell m}^{\mathrm{(ext)}}$, $A_{\ell m}$, $B_{\ell m}$ and $A_{\ell m}^{\mathrm{(in)}}$ may be found from the junction conditions, continuity of the potential $\Psi$ (in the absence of free currents) and continuity of the normal component of the magnetic field $\bm{B}=\mu_{0}(\bm{H}+\bm{M})$ on spheres $r=r_{-}$ and $r=r_{+}$. We shall consider two separate subcases:
@@
    Magnetization is of the form $\bm{M}=-\nabla\sum_{\ell,m}M_{\ell m}r^{-(\ell+1)}\mathrm{Y}_{\ell m}$. Solution of junction conditions is
@@
    Magnetization is of the form $\bm{M}=-\nabla\sum_{\ell,m}M_{\ell m}r^{\ell}\mathrm{Y}_{\ell m}$. Solution of junction conditions is
@@
  • [KDG+23] D. Kim, C. Duran, D. Giardini, A.-C. Plesa, S. C. Stähler, C. Boehm, V. Lekić, S. M. McLennan, S. Ceylan, J. F. Clinton, P. Davis, A. Khan, B. Knapmeyer-Endrun, M. P. Panning, M. Wieczorek, P. Lognonné, and W. B. Banerdt, Global Crustal Thickness Revealed by Surface Waves Orbiting Mars, Geophys. Res. Lett. 50 (2023), e2023GL103482.
  • [LL01] E. H. Lieb and M. Loss, Analysis, 2nd ed., American Mathematical Society, Providence, 2001.
@@
  • [MHM+18] J.-L. Margot, S. A. Hauck, E. Mazarico, S. Padovan, and S. J. Peale, Mercury’s Internal Structure, Mercury. The View after MESSENGER (S. C. Solomon, L. R. Nittler, and B. J. Anderson, eds.), Cambridge University Press, 2018, pp. 85–113.
  • [MJ22] A. Mittelholz and C. L. Johnson, The Martian Crustal Magnetic Field, Front. Astron. Space Sci. 9 (2022), 1–19.